OpenStudy (anonymous):

y = 15^(8x-x^2) what is the derivative?

OpenStudy (bahrom7893):

ln(y) = (8x-x^2)*ln5

OpenStudy (bahrom7893):

differentiate both sides implicitly: (1/y)(dy/dx) = (ln5)*(8-2x)

OpenStudy (bahrom7893):

Multiply both sides by y: dy/dx = (ln15)*(8-2x)*y.. by the way i had typos, both of those ln5s above should be ln15

OpenStudy (bahrom7893):

dy/dx = ln(15)(8-2x)(15^(8x-x^2))

OpenStudy (anonymous):

Thanks man you're the best!

OpenStudy (bahrom7893):

np